The length of a certain rectangle is 3 meters more than twice the width. What is the width of the rectangle if the perimeter of the rectangle is 90 meters?
Thanks,
Pat
: The length of a certain rectangle is 3 meters more than twice the width. What is the width of the rectangle if the perimeter of the rectangle is 90 meters?
Thanks,
Pat

Answer by mangopeeler07(448)

width=w
length=2w+3

2(w)+2(2w+3)=90

Divide both sides by 2
(w)+(2w+3)=45

Combine like terms
3w+3=45

Subtract 3 from both sides
3w=42

Divide both sides by 3
w=14

width=w
length=2w+3

width=14
length=2(14)+3
length=28+3
length=31

Answer:
*width=14 meters*
length=31 meters